# Deciphering Gas Dynamics and Star Formation in a z=1.1 Main Sequence Spiral Galaxy with ALMA and JWST

### v1

@@ 5 Summary and Future Prospect
This paper presents a joint analysis of a spiral galaxy, ASPECS-LP.3mm.06, utilizing JWST NIRCam WFSS and ALMA observations. By combining these two datasets, we were able to study the molecular gas dynamics and spatially resolved star formation on kiloparsec scales within this gas-rich, main-sequence galaxy. ASPECS-LP.3mm.06 was robustly detected in both the emission line and CO(2-1) emission line. The emission reveals an extended and clumpy structure, while the CO(2-1) emission indicates an extended gas disk. It exhibits a gas-rich nature with Mmol = and a moderate SFR with SFRPaα = 28.0 1.0 yr-1. The size measurements of the galaxy reveal an extended structure in F090W (5.29 0.27 kpc), F444W (4.15 0.18 kpc), and the CO integrated map (6.23 0.25 kpc). The F444W band, corresponding to the stellar mass distribution, has a slightly smaller half-light radius, which may be explained by the presence of a bulge. We modeled the kinematics of ASPECS-LP.3mm.06 and found a rotating disk with a flat rotation curve and regular kinematic features, classifying it as a dynamically cold disk galaxy. We developed two methods to enable direct comparison of NIRCam WFSS and ALMA observations. Based on this, we calculated the spatially resolved SFE within the galaxy and found a smooth SFE across both the bulge and disk regions. Despite the high gas fraction of 0.44 and high metallicity, ASPECS-LP.3mm.06 is not very dusty, with APaα = 0.13 and AV = 0.8, distinguishing it from common sub-millimeter dusty starburst galaxies. The relatively low dust content, combined with the presence of a modest bulge component (with a bulge-to-disk ratio of 0.130.01), suggests that star formation in this galaxy may be regulated by central gas turbulence. The relatively low SFR of ASPECS-LP.3mm.06, compared to galaxies with similar stellar mass and gas fraction in the literature, further supports this hypothesis. Our observations suggest that ASPECS-LP.3mm.06 is a prototypical disk galaxy at z1 and that such galaxies, without any dramatic activities like major mergers, may be undergoing an evolutionary path characterized by gradual bulge growth and self-regulated star formation. This indicates that internal processes, rather than external interactions, play a significant role in their evolution toward quiescence.

### v2

@@ 5 Summary and Future Prospect
This paper presents a joint analysis of a spiral galaxy, ASPECS-LP.3mm.06, utilizing JWST NIRCam WFSS and ALMA observations. By combining these two datasets, we were able to study the molecular gas dynamics and spatially resolved star formation on kiloparsec scales within this gas-rich, main-sequence galaxy. ASPECS-LP.3mm.06 was robustly detected in both the emission line and CO(2-1) emission line. The emission reveals an extended and clumpy structure, while the CO(2-1) emission indicates an extended gas disk. It exhibits a gas-rich nature with Mmol = and a moderate SFR with SFRPaα = 28.0 1.0 yr-1. The size measurements of the galaxy reveal an extended structure in F090W (5.29 0.27 kpc), F444W (4.15 0.18 kpc), and the CO integrated map (5.72 0.28 kpc). The F444W band, corresponding to the stellar mass distribution, has a slightly smaller half-light radius, which may be explained by the presence of a bulge. We modeled the kinematics of ASPECS-LP.3mm.06 and found a rotating disk with a flat rotation curve and regular kinematic features, classifying it as a dynamically cold disk galaxy. We developed two methods to enable direct comparison of NIRCam WFSS and ALMA observations. Based on this, we calculated the spatially resolved SFE within the galaxy and found a smooth SFE across both the bulge and disk regions. Despite the high gas fraction of 0.44 and high metallicity, ASPECS-LP.3mm.06 is not very dusty, with APaα = 0.13 and AV = 0.8, distinguishing it from common sub-millimeter dusty starburst galaxies. The relatively low dust content, combined with the presence of a modest bulge component (with a bulge-to-disk ratio of 0.130.01), suggests that star formation in this galaxy may be regulated by central gas turbulence. The relatively low SFR of ASPECS-LP.3mm.06, compared to galaxies with similar stellar mass and gas fraction in the literature, further supports this hypothesis. Our observations suggest that ASPECS-LP.3mm.06 is a prototypical disk galaxy at z1, with no clear evidence of recent major merger activity. The galaxy’s current state may reflect an evolutionary path characterized by gradual bulge growth and self-regulated star formation. This indicates that internal processes, rather than external interactions, play a significant role in their evolution toward quiescence.
@@
We sincerely thank the anonymous referee for the insightful comments, which have greatly enhanced the quality of this manuscript. We acknowledge the teams of the JWST observation programs #1180, #1895, and the HST observation program #14227 for their hard work in designing and planning these programs, and for generously making their data publicly available. ZL would like to thank Dr. Jorge González-López for discussions regarding data reduction and Ryota Ikeda for discussions on size measurement. The data presented in this paper were retrieved from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute, the specific observations analyzed can be accessed via http://dx.doi.org/10.17909/z0sb-mk09 (catalog DOI: 10.17909/z0sb-mk09). This paper makes use of the following ALMA data: ADS/JAO.ALMA#2018.1.01521.S ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. Data analysis was carried out on the Multi-wavelength Data Analysis System operated by the Astronomy Data Center (ADC), National Astronomical Observatory of Japan. ZL is supported by KAKENHI Grant No. 24KJ0394 from the Japan Society for the Promotion of Science (JSPS) and by the ALMA Japan Research Grant of the NAOJ ALMA Project, NAOJ-ALMA-342. This work is also supported by JSPS KAKENHI Grants No. 24H00002 (Specially Promoted Research by T. Kodama et al.) and No. 22K21349 (International Leading Research by S. Miyazaki et al.).
, BBarolo (Di Teodoro & Fraternali, 2015))